OpenStudy (anonymous):

simplify completely 3x+18/18

OpenStudy (lilg132):

3x + 1

OpenStudy (anonymous):

its either 3x or x

OpenStudy (lilg132):

my bad x = -1/3

OpenStudy (anonymous):

You can factor it to 3(x+6)/3(6), then take off the 3's, so you get x+6/6. I think that's as far as you can simplify it....

OpenStudy (lilg132):

because you have 3x + 18/18 18/18 = 1 which gives you 3x + 1 = 0 3x = -1 3x / 3 = -1 / 3 x = -1/3

OpenStudy (lilg132):

is my answer correct?

OpenStudy (anonymous):

the anser choices are x, 3x, x+6/6, or x+18/6

OpenStudy (anonymous):

Not if the question looks like this: \[(3x+18)\div18\]

OpenStudy (anonymous):

it's the 3rd answer- x+6/6

OpenStudy (anonymous):

you can't simplify things that are added or subtracted to x. only if it's divided or multiplied.

OpenStudy (anonymous):

it was